Check whether the sequence (an), where


an = 1/ (n+1) + 1/(n+2) +....+1/(2n) is convergent or not


1
Expert's answer
2022-06-06T14:35:11-0400

Consider


"1+\dfrac{1}{2}+...+\dfrac{1}{n}+\dfrac{1}{n+1}+...+\dfrac{1}{2n}"

"=(1+\dfrac{1}{2}+...+\dfrac{1}{n})+(\dfrac{1}{n+1}+...+\dfrac{1}{2n})"

Then


"\dfrac{1}{n+1}+...+\dfrac{1}{2n}=(1+\dfrac{1}{2}+\dfrac{1}{3}+...+\dfrac{1}{n}"

"+\dfrac{1}{n+1}+...+\dfrac{1}{2n-1}+\dfrac{1}{2n})-2(\dfrac{1}{2}+\dfrac{1}{4}+...+\dfrac{1}{2n})"

"=1-\dfrac{1}{2}+\dfrac{1}{3}-\dfrac{1}{4}+...+\dfrac{1}{2n-1}-\dfrac{1}{2n}"

"=\displaystyle\sum_{k=1}^{2n}\dfrac{(-1)^{k+1}}{k}"

Consider "\ln(1+x)" for "x\in(-1, x]"


"\ln(1+x)=\displaystyle\sum_{k=1}^{\infin}\dfrac{(-1)^{k+1}x^k}{k}"

Then for "x=1"


"\ln2=\ln(1+1)=\displaystyle\sum_{k=1}^{\infin}\dfrac{(-1)^{k+1}(1)^k}{k}"

"=\displaystyle\sum_{k=1}^{\infin}\dfrac{(-1)^{k+1}}{k}"

We see that


"\lim\limits_{n\to\infin}a_n=\lim\limits_{n\to\infin}(\dfrac{1}{n+1}+...+\dfrac{1}{2n})"

"=\lim\limits_{n\to\infin}\displaystyle\sum_{k=1}^{2n}\dfrac{(-1)^{k+1}}{k}=\displaystyle\sum_{k=1}^{\infin}\dfrac{(-1)^{k+1}}{k}=\ln2"

Therefore he sequence "(a_n)" is convergent.